# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 3.2.3. Approximation of the problem
 $\text{ either }\frac{w_{j}^{\varepsilon}(x+re_{1})-w_{j}^{\varepsilon}(x)}{r}\in\omega_{\varepsilon}\quad\text{ or }\quad\frac{w_{j}^{\varepsilon}(x+re_{1})-w_{j}^{\varepsilon}(x)}{r}\in\mathbb{R}^{2}\setminus\omega_{\varepsilon}.$ (3.25)
Let us characterize $\omega_{\varepsilon}$. If we consider, for any $\varepsilon\in[0,1]$, the continuous map defined in complex notation by:
$\Phi_{\varepsilon}(z)=ze_{1}+(1-\varepsilon)h(z)\bar{e}_{1}=z+(1-\varepsilon)h(z),$
then $\Phi_{\varepsilon}$ is injective from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$, thanks to (3.1). Thus, it is a homeomorphism onto its open image. The image is also closed thanks to the boundedness of $h$ (3.2), and hence $\Phi_{\varepsilon}(\mathbb{R}^{2})=\mathbb{R}^{2}$, $\forall\varepsilon\in[0,1]$. Noticing that $\Phi_{\varepsilon}(\gamma(t))=\Gamma_{\varepsilon}(t)e_{1}$ for all $t\in[0,1]$ by (2.2) and recalling (3.18), we deduce:
$\omega_{\varepsilon}=\Phi_{\varepsilon}(B_{\tau}(a)).$ (3.26)
Letting $\varepsilon\to 0$ in (3.25), since $\Phi_{\varepsilon}$ converges uniformly to $\Phi_{0}$ on $\mathbb{R}^{2}$ we find that for all $r<\delta$:
 $\text{ either }\frac{w_{j}(x+re_{1})-w_{j}(x)}{r}\in\overline{\omega_{0}}\quad\text{ or }\quad\frac{w_{j}(x+re_{1})-w_{j}(x)}{r}\in{\mathbb{R}^{2}\setminus\omega_{0}},\quad\text{for all $x\in B_{\delta}$.}$ (3.27)
If, by contradiction, the first alternative occurred for infinitely many $j$, then we would find
$Dw_{j}(x)e_{1}=\Phi_{0}(\partial_{z}w(x))\in\overline{\omega_{0}}=\Phi_{0}(\overline{B_{\tau}(a)}),\text{ for a.e. }x\in B_{\delta}.$
In turn, since $\Phi_{0}$ is a homeomorphism, we deduce that $\partial_{z}w_{j}(x)\in\overline{B_{\tau}(a)},\text{ for a.e. }x\in B_{\delta}$ and infinitely many $j\geq J$. By Theorem 1.1, this would also imply that
 $\text{ $\partial_{z}w_{\infty}(h)\in\overline{B_{\tau}(a)}\subset(\overline{C})^{c}$ a.e. in $B_{\delta}$},$ (3.28)
which is against our assumption $\partial_{z}w_{\infty}](B_{r})\subset C$. Therefore, the second alternative in (3.27) must occur. Following the same reasoning of the previous contradiction argument, this leads to
$|\partial_{z}w_{j}(x)-a|\geq\tau,\quad\text{ for a.e. $x\in B_{\delta}$},\forall j\geq J,$
which is precisely the conclusion of Theorem 3.2.∎
4. Theorem B: the size of non-differentiability points
In this section we give the proof of Theorem B, that we deduce from the following result valid in $\mathbb{R}^{n}$.
Proposition 4.1.
@@ 4. Theorem B: the size of non-differentiability points
see [21, Definition 2.1]. Our definition differs from that of [21, Definition 2.1] by a constant factor, which is anyway irrelevant for what we need to show. We claim that:
@@
The $\sigma$-subadditivity of $\mathcal{H}^{n-1}$, [21, Theorem 2.1, Claim # 2], and (4.5) would then conclude the proof. We are only left to show (4.6). To this aim, pick any $x\in E_{p,r}$. Then, monotonicity and convexity of $\sigma$ imply:
@@
As $r$ is fixed and $E_{p,r}\subset B_{1}$, we can cover $E_{p,r}$ with finitely many balls $\overline{B_{r}(x_{i})}$ centered at $\{x_{i}\}_{i=1}^{N}\subset E_{p,r}$. From Vitali’s Covering Theorem, [21, Theorem 1.24], we find a subset of centers $x_{i_{1}},\dots,x_{i_{M}}$ such that
@@
4.1. Proof of Theorem B
Let $w$ solve (1.8). Writing $w=(u,v)$, by Proposition 2.2 we can find a monotone field $G:[Du]\to\mathbb{R}^{2}$ fulfilling (4.1) such that $u$ solves (1.1). To show Theorem B we prove that $S^{c}\subset\operatorname{Reg}(w).$ Let then $x_{0}\in S^{c}$. This means that there exists a sequence $r_{n}\to 0$ such that
$\lim_{n}\fint_{B_{r_{n}}(x_{0})}\sigma\left(\left|Du(y)-\fint_{B_{r_{n}}(x_{0})}Du(z)dz\right|\right)dy=0.$ (4.12)
We may assume, up to a non-relabeled subsequence, that $\fint_{B_{r_{n}}(x_{0})}Du\to a\in\mathbb{R}^{2}.$ Up to considering another non-relabeled subsequence, we can assume that $w_{r_{n},x_{0}}$, recall (1.13), converges locally uniformly and, thanks to Theorem 1.1, strongly in $W^{1,1}_{\operatorname{loc}}$ to $w_{\infty}=(u_{\infty},v_{\infty})$. Now the strong convergence and (4.12) imply that $Du_{\infty}\equiv a$ on $B_{1}$. From (1.9)-(1.6)-(1.14), we infer that $Dw_{\infty}\equiv A$ in $B_{1}$, where $A\in K$ is the only matrix in $K$ whose first row is $a$. Hence $w_{\infty}$ is differentiable at 0, and $x_{0}\in\operatorname{Reg}(w)$.
5. Theorem C: Monge-Ampère measure associated to solutions of elliptic PDEs
This section is devoted to showing Theorem C. For $u,v\in W^{1,2}(B_{1})$, we define the distribution
@@ 5. Theorem C: Monge-Ampère measure associated to solutions of elliptic PDEs
where $J$ is defined in (2.6). Recalling (2.1), if $u,v\in W^{2,2}$, a direct computation shows that $\mathcal{D}(u,v)=\frac{1}{2}\langle D^{2}u,\operatorname{cof}D^{2}v\rangle$, so that $\mathcal{D}(u,u)=\det(D^{2}u)$. Theorem C follows from Theorem 5.1 and Corollary 5.
@@
which would conclude the proof. Let us show the claim. By Proposition 2.2 we have that $u$ is a Lipschitz solution to (1.1) with $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ satisfying (1.5). By [35, Lemma A.4], $u$ is a strong $H^{1}$ limit of smooth functions $u_{\varepsilon}$ solving
@@
where $G^{\varepsilon}$ are smooth fields satisfying (1.5) with $\sigma(t)=c_{\varepsilon}t^{2}$ for some $c_{\varepsilon}>0$, hence $DG^{\varepsilon}+(DG^{\varepsilon})^{T}\geq 2c_{\varepsilon}$. This and (5.3) imply that $\mathcal{D}(u^{\varepsilon},u^{\varepsilon})=\det(D^{2}u^{\varepsilon})\leq 0$ in $\mathcal{D}^{\prime}(B_{1})$. By the strong $H^{1}$ convergence provided by [35, Lemma A.4], we deduce the same for $\mathcal{D}(u,u)$. ∎
@@
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (1.9), and let $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ be a solution to (1.8). Consider the matrix-valued measure $\mathcal{D}(w)$ defined as in (5.2). Factorize it as a $\mathcal{D}(w)=P\mu$, where $\mu$ is a finite, positive measure on $B_{1}$, and $P\in L^{\infty}(B_{1},\operatorname{Sym}^{-}(2);\mu)$. Then, for any $w_{\infty}\in\mathcal{B}(w)(x_{0})$, recall (1.14),
@@
6. Theorem D: Regularity of inclusions into curves
In this section we show Theorem D. The proof is quite long, so it is useful to split it into steps. Throughout, we denote by $\Gamma$ the $C^{1}$ curve fulfilling (1.9) for which, as in Theorem D, $w$ solves
$Dw\in\Gamma\text{ a.e. in $\Omega$}.$ (6.1)
As some results depend on the global geometry of the domain, it is convenient to consider a general open, connected, simply connected set $\Omega$ in (6.1). We let $\Gamma=\gamma(I)$ for some $I=[a,b]$ or $\mathbb{R}/L{\mathbb{Z}}$ and $\gamma\in C^{1}(I,\mathbb{R}^{2\times 2})$ a homeomorphism onto $\Gamma$ with $|\gamma^{\prime}|>0$ on $I$. The proof will be split into steps:
@@ 6. Theorem D: Regularity of inclusions into curves
  • The fifth and sixth step, contained in §6.5-6.6, are devoted to show that $w$ is an entropy solution to $Dw\in\Gamma$ and to deduce from this the structure of the singular set, respectively.
Theorem D is then a direct consequence of Corollary 6.1, Lemma 6.2, Propositions 6.5.1-6.5.2 and Lemma 6.6.
@@ 6.1. Step 1: Viscosity properties
This proof is essentially contained in [54, Lemma 3]. Indeed we have:
@@
Thanks to the previous reductions, we now fix $\zeta\in C^{1}(B_{1})$ such that $u-\zeta$ has a strict minimum at 0, and we let $K_{1}=\pi_{1}(K)\subset\mathbb{R}^{2}$, as in Proposition 2.2. To show $D\zeta(x_{0})\in K_{1}$, we follow the strategy of [18, §4.2]. For small $\delta>0$, we consider the open set $\Omega_{\delta}\subset\subset\Omega$ given by the connected component of $\{u-\zeta<\delta\}$ which contains 0. Denote by $\nu_{\delta}$ and $\mu_{\delta}$ the following probability measures:

### v2

@@ 4.1.3. Approximation of the problem
or $\frac{w^{\varepsilon}(x+te_{1})-w^{\varepsilon}(x)}{t}\in\Phi^{\varepsilon}(\mathbb{R}^{2}\setminus\overline{B_{R}})\quad\forall(x,t)\in B_{\eta}\times(0,\eta)\,.$
One of these must be satisfied for infinitely many $\varepsilon\in(0,\varepsilon_{0})$, and from the uniform convergences of $w^{\varepsilon}$ to $w$ and $\Phi^{\varepsilon}$ to $\Phi^{0}\colon z\mapsto g(z)e_{1}$ we deduce
either $\frac{w(x+te_{1})-w(x)}{t}\in\Phi^{0}(\overline{C})\quad\forall(x,t)\in B_{\eta}\times(0,\eta)\,$,
or $\frac{w(x+te_{1})-w(x)}{t}\in\Phi^{0}(\mathbb{R}^{2}\setminus{B_{R}})\quad\forall(x,t)\in B_{\eta}\times(0,\eta)\,.$
Letting $t\to 0$, at every differentiability point $x\in B_{\eta}$ of $w$ we have therefore
$\Phi^{0}(\partial_{z}w(x))=Dw(x)e_{1}\in\Phi^{0}(\overline{C})\cup\Phi^{0}(\mathbb{R}^{2}\setminus B_{R})\,$,
hence $\partial_{z}w(x)\in\overline{C}\cup(\mathbb{R}^{2}\setminus B_{R})$, as $\Phi^{0}$ is a homeomorphism. Combining this with the fact that $Dw\in K\subset g(B_{R})$ a.e., which implies $\partial_{z}w\in B_{R}$ a.e., we eventually deduce that $\partial_{z}w(x)\in\overline{C}$ for a.e. $x\in B_{\eta}$. Setting $\delta=\min(\alpha,\eta)>0$ concludes the proof of Theorem 4.3. ∎
5. The size of non-differentiability points
In this section we show the following result valid in $\mathbb{R}^{n}$.
Theorem 5.1.
@@ 5. The size of non-differentiability points
see [23, Definition 2.1]. Our definition differs from that of [23, Definition 2.1] by a constant factor, which is anyway irrelevant for what we need to show. We claim that:
@@
The $\sigma$-subadditivity of $\mathcal{H}^{n-1}$, [23, Theorem 2.1, Claim # 2], and (5.5) would then conclude the proof. We are only left to show (5.6). To this aim, pick any $x\in E_{p,r}$. Then, monotonicity and convexity of $\sigma$ imply:
@@
As $r$ is fixed and $E_{p,r}\subset B_{1}$, we can cover $E_{p,r}$ with finitely many balls $\overline{B_{r}(x_{i})}$ centered at $\{x_{i}\}_{i=1}^{N}\subset E_{p,r}$. From Vitali’s Covering Theorem, [23, Theorem 1.24], we find a subset of centers $x_{i_{1}},\dots,x_{i_{M}}$ such that
@@
6. Proof of Theorem 1.1
Before showing the Theorem, let us introduce some useful notation. Given a map $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ and any point $x_{0}\in B_{1}$, we set:
$w_{r,x_{0}}(h)\doteq\frac{w(x_{0}+rh)-w(x_{0})}{r}.$ (6.1)
As $w$ is Lipschitz, this sequence of rescalings is precompact in $C^{0}(B)$, for any ball $B\subset\mathbb{R}^{2}$. Therefore, we can introduce the set $\mathcal{B}(w)(x_{0})$ of blowups of $w$ at $x_{0}$ to be the collection of maps obtained as locally uniform limits of any subsequence extracted from $w_{r,x_{0}}$. Observe that, by Rademacher’s theorem, $w$ is differentiable at a.e. $x_{0}\in B_{1}$ and thus $\mathcal{B}(w)(x_{0})=\{L_{Dw(x_{0})}\}$ for any such $x_{0}$. Given now a solution $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2}),B_{1}\subset\mathbb{R}^{2}$ to (2.3) with $K$ enjoying property (2.4), we further observe that:
[…]
due to the strong $W^{1,1}_{\operatorname{loc}}$ convergence of the gradients of rescalings provided by Theorem 2.2. Through blowups, we define the regular and the singular sets of $w$:
Definition 6.1.
For $w\in\operatorname{Lip}(\Omega,\mathbb{R}^{2})$, $\Omega\subset\mathbb{R}^{2}$ open, we let
[…]
Combining now Theorems 4.1-5.1, we get the following precise version of Theorem 1.1, restated in terms of differential inclusion thanks to Propositions 3.2-3.2.
Theorem 6.2.
Let $K\subset\mathbb{R}^{2\times 2}$ be compact and fulfill (2.4) and let $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$, $B_{1}\subset\mathbb{R}^{2}$, solve (2.3). Then, $w$ is differentiable over $\operatorname{Reg}(w)$, $Dw|_{\operatorname{Reg}(w)}$ is continuous, and $\mathcal{H}^{1}(\operatorname{Sing}(w))=0$.
Proof.
Let $x_{0}\in\operatorname{Reg}(w)$. By definition, there exists $v\in\mathcal{B}(w)(x_{0})$ which is differentiable at 0. By (6.2), $Dv\in K$ a.e. in $\mathbb{R}^{2}$. Since $v$ is differentiable at 0, Theorem 4.1 applied to the family $v_{r,0}$, compare (6.1), implies that $\operatorname{diam}([Dv](0))=0$. Now, as $v\in\mathcal{B}(w)(x_{0})$, then $v$ is the local uniform limit in $\mathbb{R}^{2}$ of $(w_{r_{j},x_{0}})_{j\in{\mathbb{N}}}$, and we can employ the same reasoning to deduce from Theorem 4.1 that $\operatorname{diam}([Dw](x_{0}))=0$. Therefore, it consists of a single matrix, and differentiability and continuity at that point follow directly from the definition (2.7). To conclude the proof, it suffices to show that $\operatorname{Sing}(w)\subset S$, where $S$ is the $\mathcal{H}^{1}$-negligible set provided by Theorem 5.1. Here, Theorem 5.1 is applied to the first component $u$ of $w=(u,v)$, which solves (1.1) for the strictly monotone field $G$ associated to $K$ by Proposition 3.2. Equivalently, we show that $S^{c}\subset\operatorname{Reg}(w)$. Let then $x_{0}\in S^{c}$. This means that there exists a sequence $r_{n}\to 0$ such that
$\lim_{n\to+\infty}\fint_{B_{r_{n}}(x_{0})}\sigma\left(\left|Du(y)-\fint_{B_{r_{n}}(x_{0})}Du(z)dz\right|\right)dy=0.$ (6.3)
We may assume, up to a non-relabeled subsequence, that $\fint_{B_{r_{n}}(x_{0})}Du\to a\in\mathbb{R}^{2}.$ Up to considering another non-relabeled subsequence, we can assume that $w_{r_{n},x_{0}}$, recall (6.1), converges locally uniformly and, thanks to Theorem 2.2, strongly in $W^{1,1}_{\operatorname{loc}}$ to $w_{\infty}=(u_{\infty},v_{\infty})$. Now the strong convergence and (6.3) imply that $Du_{\infty}\equiv a$ on $B_{1}$. From (2.4)-(2.2)-(6.2), we infer that $Dw_{\infty}\equiv A$ in $B_{1}$, where $A\in K$ is the only matrix in $K$ whose first row is $a$. Hence $w_{\infty}$ is differentiable at 0, and $x_{0}\in\operatorname{Reg}(w)$. ∎
7. Monge-Ampère measure associated to solutions of elliptic PDEs
For $u,v\in W^{1,2}(B_{1})$, we define the distribution
@@ 7. Monge-Ampère measure associated to solutions of elliptic PDEs
where $J$ is defined in (3.6). Recalling (3.1), if $u,v\in W^{2,2}$, a direct computation shows that $\mathcal{D}(u,v)=\frac{1}{2}\langle D^{2}u,\operatorname{cof}D^{2}v\rangle$, so that $\mathcal{D}(u,u)=\det(D^{2}u)$. In the literature, $\mathcal{D}(u,u)$ is usually called the very weak Hessian Jacobian of $u$. The main result of this section is the following:
@@
which would conclude the proof. Let us show the claim. By Proposition 3.2 we have that $u$ is a Lipschitz solution to (1.1) with $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ satisfying (2.1). By [37, Lemma A.4], $u$ is a strong $H^{1}$ limit of smooth functions $u_{\varepsilon}$ solving
@@
where $G^{\varepsilon}$ are smooth fields satisfying (2.1) with $\sigma(t)=c_{\varepsilon}t^{2}$ for some $c_{\varepsilon}>0$, hence $DG^{\varepsilon}+(DG^{\varepsilon})^{T}\geq 2c_{\varepsilon}$. This and (7.3) imply that $\mathcal{D}(u^{\varepsilon},u^{\varepsilon})=\det(D^{2}u^{\varepsilon})\leq 0$ in $\mathcal{D}^{\prime}(B_{1})$. By the strong $H^{1}$ convergence provided by [37, Lemma A.4], we deduce the same for $\mathcal{D}(u,u)$. ∎
@@
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (2.4), and let $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ be a solution to (2.3). Consider the matrix-valued measure $\mathcal{D}(w)$ defined as in (7.2). Factorize it as $\mathcal{D}(w)=P\mu$, where $\mu$ is a finite, positive measure on $B_{1}$, and $P\in L^{\infty}(B_{1},\operatorname{Sym}^{-}(2);\mu)$. Then, for any $w_{\infty}\in\mathcal{B}(w)(x_{0})$, recall (6.2),
@@
Let us end this section with a few general comments. Firstly, similar properties hold for infinity harmonic functions [35, Theorem 1.5]. Secondly, these results suggest the following conjectural analysis of blowups: one could hope to deduce from Corollary 7 that $w_{\infty}\in\mathcal{B}(w)(x_{0})$ is linear at points where $\mu(\{x_{0}\})=0$ and that $w_{\infty}$ is one-homogeneous at points where $\mu(\{x_{0}\})\neq 0$. This would tell us that $\operatorname{Sing}(w)$ is actually countable, since a finite measure can have at most countably many atoms. The reason for this hope is that the equation $\det(D^{2}u)=0$ in an open set $\Omega$ implies a very rigid structure of $Du$: essentially, that $Du$ is constant along segments connecting the boundary. This, in turn, yields that $Du$ is constant if $u$ is defined in $\mathbb{R}^{2}$. Such rigidity property has been shown assuming $u\in C^{2}$ [28], $u\in W^{2,\infty}$ [33, §2.6], $u\in W^{2,2}$ [51] and $u$ in the class MA [31]. In all of these cases, $Du$ is (at least weakly) differentiable. Here, however, the only way we can interpret the distribution $\det(D^{2}u)$ is in the very weak sense, see the distribution $\mathcal{D}(u,u)$ in (7.1). The equation $\mathcal{D}(u,u)=0$ is extremely flexible, even if $u\in C^{1,\alpha}$ for $\alpha>0$, see [40, 7, 6]. In a sense, a function $u$ which satisfies $\mathcal{D}(u,u)=0$ and also (1.1) for a strictly monotone $G$, lies in a realm between the known rigidity and flexibility statements. The fact that some rigidity can be expected can be seen in Proposition 9.2, where Corollary 7 will be instrumental to exclude some types of blowups.
8. Regularity of inclusions into curves
In this section we study inclusions into elliptic curves. Namely, we assume that $K=\Gamma\subset\mathbb{R}^{2\times 2}$ satisfies (2.4) and is a $C^{1}$ curve, that is, a connected, compact $C^{1}$ submanifold of $\mathbb{R}^{2\times 2}$ of dimension one, which may have a nonempty boundary. Hence $\Gamma=\gamma(I)$ for some $I=[a,b]$ or $\mathbb{R}/L{\mathbb{Z}}$ and $\gamma\in C^{1}(I,\mathbb{R}^{2\times 2})$ is a homeomorphism onto $\Gamma$ with $|\gamma^{\prime}|>0$ on $I$. As some results depend on the global geometry of the domain, it is convenient to consider a general open, connected, simply connected set $\Omega$ in (8.1). As previously mentioned, the next Theorem generalizes the main result of [39], where the same conclusion was obtained under the assumption that (2.4) is satisfied with $\sigma(t)=ct^{4}$ for some $c>0$.
Theorem 8.1.
Let $\Omega\subset\mathbb{R}^{2}$ be open. If $w\in\operatorname{Lip}(\Omega,\mathbb{R}^{2})$ satisfies
$Dw\in\Gamma\text{ a.e. in $\Omega$},$ (8.1)
then $\operatorname{Sing}(w)$ is locally finite. If $\Omega$ is convex, then $\operatorname{Sing}(w)$ contains at most two points. Moreover, $\operatorname{Sing}(w)$ is empty if one of the following sufficient conditions is satisfied.
  • $\Gamma$ is simply connected, that is, $I=[a,b]$;
In fact, under the last assumption, the conclusion holds even if $\Gamma$ is not assumed $C^{1}$, see Corollary 8.1. Aside from its independent interest, this result will be crucial for the proof of Theorem 9.1. As the proof is quite long, it is useful to split it into steps:
@@ 8. Regularity of inclusions into curves
  • The fifth and sixth step, contained in §8.5-8.6, are devoted to show that $w$ is a zero-entropy solution to $Dw\in\Gamma$ and to deduce from this the structure of the singular set, respectively.
Theorem 8.1 is then a consequence of Corollary 8.1, Lemma 8.2, Propositions 8.5.1-8.5.2 and Lemma 8.6.
@@ 8.1. Step 1: Viscosity properties
This proof is essentially contained in [56, Lemma 3]. Indeed we have:
@@
Thanks to the previous reductions, we now fix $\zeta\in C^{1}(B_{1})$ such that $u-\zeta$ has a strict minimum at 0, and we let $K_{1}=\pi_{1}(K)\subset\mathbb{R}^{2}$, as in Proposition 3.2. To show $D\zeta(x_{0})\in K_{1}$, we follow the strategy of [20, §4.2]. For small $\delta>0$, we consider the open set $\Omega_{\delta}\subset\subset\Omega$ given by the connected component of $\{u-\zeta<\delta\}$ which contains 0. Denote by $\nu_{\delta}$ and $\mu_{\delta}$ the following probability measures: